OpenStudy (anonymous):

Evaluate cube root of 5 multiplied by square root of 5 over cube root of 5 to the power of 5 @pgpilot326

OpenStudy (anonymous):

it looks like this @pgpilot326

OpenStudy (anonymous):

you can rewrite using fractional exponents and then simplify using the properties of exponents.

OpenStudy (anonymous):

this confuses me my teachers did not explain this

OpenStudy (anonymous):

hold on a sec...

OpenStudy (anonymous):

http://www.mathsisfun.com/algebra/exponent-laws.html check this out and then get back to me if you have any questions.

OpenStudy (anonymous):

@pgpilot326 im confused with the exponent in the front what do i do with that is it the same as it being behind the number

OpenStudy (anonymous):

please write an example because i'm not undersatnding what you're saying

OpenStudy (anonymous):

y =^3 sqrt x + 2

OpenStudy (anonymous):

please use the equation editor...

OpenStudy (anonymous):

\[y =^3 \sqrt x+2\]

OpenStudy (anonymous):

\[\sqrt[3]{x} = x^{\frac{1}{3}}\]in general, \[\sqrt[n]{x^m}=\left( \sqrt[n]{x} \right)^m=x^{\frac{n}{m}}\]

OpenStudy (anonymous):

did you look at the link?